OpenStudy (anonymous):

write an equation in standard form using only intergers for the line descrived. The line perpendicular to x-3y=7 and containing (-2,5)

OpenStudy (anonymous):

do you know how to find the slope of the line x-3y=7 ?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

perpendicular lines have negative reciprocal slope with the slope of the original line. So first thing you have to do is find the slope of the given line.

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

right. then get the negative reciprocal of 1/3. which will be..?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

great :) so so far you have the slope of the perpendicular line which is m=-3 \((y-y_1)= m(x-x_1)\) are you familiar with this?

OpenStudy (anonymous):

or y=mx+b

OpenStudy (anonymous):

yeah same thing. but \((y-y_1)= m(x-x_1)\) is easier when doing the standard form. where \((x_1 , y_1)\) is your given point. substitute all the values given \(m=-3, x_1=-2, y_1=5\)

OpenStudy (anonymous):

3x+y=-1

OpenStudy (anonymous):

awesome! you don't need any help lol. you can do it :)

OpenStudy (anonymous):

ohh yea lol

OpenStudy (anonymous):

keep up the goodwork ;)